OpenStudy (anonymous):

4=a^-2 What is the next step? Original question: Give a possible formula for a function (-1,8) (1,2)

OpenStudy (anonymous):

5-3.x

OpenStudy (anonymous):

try linear first

OpenStudy (anonymous):

8=Qnaut a^(-1) over top of 2=Qnaut a^(1) Qnaut cancels top and bottom, and a^1 is brought up top and its a^-1(-1) on top now. 8/2 is 4 so I got 4=a^-2

OpenStudy (anonymous):

Its in the family of exponential functions

OpenStudy (anonymous):

It maybe a quotient of some sort

OpenStudy (anonymous):

This is review in the beginning of a Calc class

OpenStudy (anonymous):

The answer is y=4(1-2^-x)

OpenStudy (anonymous):

I guess so ? I'm really trying hard to get better at math So Im not sure about linear or not.

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

I have just been working with point slope form too, but I can't do it!

OpenStudy (anonymous):

oh crap hold on I forgot I could get M that way

OpenStudy (anonymous):

Im getting y=-3x+6

OpenStudy (anonymous):

8-2/-1-1 =-3 which is M correct?

OpenStudy (anonymous):

y=-3x+5 :) ok now how does this eventually equal y=4(2^-x) the answer given in the book

OpenStudy (anonymous):

what I described up top with Qnaut thats how you solve these I did some in my Math Lab at school already

OpenStudy (anonymous):

the question is give a possible formula for those two points

OpenStudy (anonymous):

just a variable like x

OpenStudy (anonymous):

ok thats where I was

OpenStudy (anonymous):

still doesnt match book answer though

OpenStudy (anonymous):

where did that go, I was trying 1/a^2 that seems like a good idea

OpenStudy (anonymous):

ok yes I can get a = +/- 1/2 now

OpenStudy (campbell_st):

isn't the function \[f(x) = \frac{1}{2^{(x - 2)}}\] that satisfies the the 2 points that you have

OpenStudy (anonymous):

the answer in the book is y=4(2^-x)

OpenStudy (campbell_st):

oh and my solution \[4a^2 = 1\] so \[a^2 = \frac{1}{4}\] \[a = \pm \frac{1}{2}\]

OpenStudy (campbell_st):

well look at 4 its 2^2 so you have \[y = 2^2(2^{-x}) \] then \[y = 2^{2 - x}\]

OpenStudy (anonymous):

where does the 1/2 come in from solving for a?

OpenStudy (campbell_st):

lol...what is the square root of 1/4 ...?

OpenStudy (anonymous):

Since the exponential formula y = Po a^x From (-1,8) (1,2) : y2/y1 = a^x2/ a^x1 = a ^ ( x2 - x1) 8/2 = a^(-2) 4 = 1/a² => a = ± 1/2 2 = Po ( 1/2)^1 => Po = 4 Now plug into: y = Po a^x = 4 (1/2)^x

OpenStudy (anonymous):

@campbell_st It takes me forever to understand what the poster need :'(

OpenStudy (anonymous):

how do you get from solving for a to y=2^2 equation?

OpenStudy (anonymous):

how do I work the 1/2 into the final equation?

OpenStudy (campbell_st):

well you said the textbook answer is \[y= 4(2^{-x})\] and \[4 = 2^2\] then using the index laws for multiplication.... add the powers it can be written as \[y = 2^{2 - x}\]

OpenStudy (campbell_st):

you don't need to work 1/2 into your answer as \[2^{-x} = \frac{1}{2^x}\]

OpenStudy (anonymous):

ok got it! man this is tough for me to see the step I need more practice

OpenStudy (campbell_st):

and thats why \[f(x) = \frac{1}{2^{x - 2}} \] fits your points

OpenStudy (campbell_st):

yep... lots of work on index laws....

OpenStudy (anonymous):

index laws ok now I know what they are called at least thanks guys!

OpenStudy (anonymous):

Actually, it's not difficult as it looks The first step is using ratio => a The second step is plug a and 1 point into the exponential formula => Po That's all :)

OpenStudy (anonymous):

You make it difficult because you didn't point out find the exponential formula!